OpenStudy (anonymous):

2√3 + 1 ___________ 5 - √3

OpenStudy (anonymous):

\[2\sqrt{3}+1 \div 5-\sqrt{3}\]

OpenStudy (anonymous):

\[1/2+\sqrt{2}/11\]